If the reverse bias voltage of a silicon varactor is increased by a factor of 2 , the
corresponding transition capacitance
1
Expert's answer
2020-09-28T08:10:33-0400

Capacitance:

"C=\frac{q}{U}"


"C_1=\frac{q}{U_1}"

"C_2=\frac{q}{U_2}"

where "U_2=2U_1" by the condition.


Divideng the equations, we get:

"\frac{C_2}{C_1}=\frac{U_1}{U_2}=\frac{U_1}{2U_1}=\frac{1}{2}"

Capacitance will decrease by 2 times.